Joan is 6 years older than his cousin Liza, and the product od their ages is 9 less than 16 times the sum of their ages. How old is each?
Answer by askmemath(368)

Ley Liza's age = a
John's age = a+6
Product of their ages = a(a+6) = 16(Sum of their ages ) -9
a^2+6a = 16(a+a+6) - 9
a^2 +6a = 16(2a+6)-9
a^2+6a = 32a+96-9 = 32a - 87
Subtracting 32a+96-9 from both sides
a^2 +6a -32a-87
a^2-26a -87 = 0
Factorising
a^2 -29a+3a-87=0
a(a-29) +3(a-29) = 0
(a+3)(a-29) = 0
Which means that a =-3 or a=29
Since age cannot be negative
a = 29
John's age = 29+6 = 35